# Efficient 3D-Aware Facial Image Editing via Attribute-Specific Prompt Learning

### v1

@@ 3.5 Efficiency
In addition to enabling the editing of novel attributes defined through language, our method is significantly more efficient than other methods. Firstly, our approach swiftly learns to edit new attributes with a small training time (4 to 8 minutes), a notable contrast to PREIM3D’s requirement of several hours of training alongside the necessity of a pre-trained attribute-specific classifier. Secondly, despite employing a language encoder, our method’s inference time is comparable to that of PREIM3D. This efficiency stems from the fixed nature of the prompt post-training, allowing us to compute prompt features once and reuse them efficiently. Finally, our method introduces a sharing mechanism for style mappers across multiple attributes, delivering benefits in both storage and computation.
@@ 4 Experiment
Implementation Details. For most of our experiments, we used a pre-trained 3D generative model GMPI [50] as our base 3D GAN and kept it frozen except for the alpha maps branch. We use a pre-trained CLIP [32] model to learn the attribute-specific prompt. CLIP’s text encoder is utilized for training and inference, and its image encoder is used only during training. To train our LAE module, we set $\lambda_{dclip}=1.0$, $\lambda_{sc}=0.8$, $\lambda_{id}=0.8$, $\lambda_{idvc}=0.5$, $\lambda_{latent}=0.5$ and $\lambda_{\alpha}=0.5$. We optimize using Adam [23] with a learning rate of $0.001$ and parameters $\beta_{1}=0.9$ and $\beta_{2}=0.95$. Similar to GMPI [50], we use $32$ planes for training and $96$ planes for inference and set near and far depth for MPI as $0.95/1.12$, use the depth normalization as [50]. To further evaluate, we integrate our method to other 3D generative model like EG3D [7], StyleNeRF [14] and CIP3D [51]. Our method is extremely efficient in terms of both computation and space requirements, as it only requires style tokens and style mappers to be trained and stored. Our code and models will be publicly released.
@@ 4.1 Quantitative Results
As GMPI [50] and EG3D [7] lacks attribute editing capabilities, we present in Table 3 a single averaged value across 1000 images. In contrast, our method calculates this value individually across three attributes (smile, makeup, and age) using 1000 images for each attribute, reporting the average value. In a comparison of the depth and pose accuracy between GMPI [50], EG3D [7] and our method. As the table shows, our method faithfully preserves depth and pose accuracy across attributes. For instance, concerning smile attribute editing, our method exhibits a depth accuracy of 0.51, marginally lower than GMPI’s 0.49 (lower is better). While using the GMPI backbone, when integrating the LAE Module with the EG3D, we can observe a similar pattern. Similar observations are evident for makeup and age attributes. These findings highlight our method’s capability to maintain 3D geometry while facilitating attribute editing.
@@ 4.2 Qualitative Results
Firstly, to demonstrate the efficacy of our method, we showcase qualitative results and compare them with the GMPI and the state-of-the-art PREIM3D [24] in Figure 3. Following the approach outlined in [24], we uniformly sampled 9 images with yaw angles ranging from -30° to 30° and pitch angles between -20° and 20°. To facilitate comparative analysis, we replicate their results by applying pre-trained weights provided by GMPI [50] and PREIM3D [24].
@@
To showcase our method’s efficacy in preserving identity and camera poses, we compare it with GMPI in Figure 5. This comparison spans two attributes observed across four randomly selected camera angles. The figure demonstrates how our method handles editing intricate and challenging attributes, such as specific emojis or blue eye colors, without compromising camera angles or altering the subject’s identity. We show the result of editing for two attributes (makeup and surprised face) at randomly selected angles in Figure 4.
Further, we use CelebA-HQ [19] to illustrate our method’s ability to edit real images. We first invert images to obtain their latent code using E4E [42]. Subsequently, these images are edited using our method, and the results of this experiment are shown in Figure 5(b). Our method maintains the identity and preserves the camera pose while successfully altering attributes. It’s essential to note that the preservation quality of our method is inherently constrained by GMPI.
@@ 4.3.2 Robustness and Biasness
To delve deeper into understanding the influence of attributes on each other, such as how changing "blue eyes" may affect "skin tone" as shown in row 2 of Figure. 5(a), we explored the bias present in the CLIP model. Our method finds semantic direction in the latent space of 3D-aware GANs via. text, encoded by CLIP. Attribute editing accuracy and control rely on CLIP’s encoding, potentially inheriting biases, noted in previous works[3]. To understand the biases of the clip, we performed a simple experiment by encoding 10 variations of prompts for white, brown, and black faces and blue eyes with CLIP and calculating the distance of the faces with blue eyes. We found that the encoding of the white face is closest to the blue eyes: white $0.035$, brown $0.051$, and black $0.065$.
4.4 Ablation Studies
Figure 8(a) highlights the impact of introducing proposed losses and the LAE Module into the 3D generation method for the attribute "curly hair". Integrating the LAE Module with $L_{LAE}$ loss brings about changes in the desired attribute, emphasizing the significance of using $L_{LAE}$ loss during the training of the LAE Module. However, despite altering the desired attributes, the LAE module falls short of preserving the identity and poses of the edited images. This issue is addressed by introducing $L_{id}+L_{idvc}$, resulting in images with desired attributes while maintaining a similar identity to the original image, though poses are still not adequately preserved. Our final approach, employing $L_{total}$, effectively preserves poses, identity, and generates 3D-aware edited images
@@ 4.4 Ablation Studies
5 Conclusion
This paper presents an efficient pipeline for editing 3D-aware and view-consistent facial image attributes specified through textual prompts. Our method comprises a text-driven Latent Attribute Editor (LAE) alongside a 3D GAN. The LAE integrates learned style tokens and a style mapper, leveraging a pre-trained CLIP model to find semantic editing directions within 3D GAN latent space. The text-driven approach and utilization of style tokens render our method exceptionally efficient in handling novel editing directions. Additionally, we employ several 3D-aware identity and pose preservation losses to ensure view consistency in the generated images. Our method’s effectiveness is validated through a comprehensive array of qualitative and quantitative experiments.

### v2

@@ 3.5 Efficiency
In addition to enabling the editing of novel attributes defined through language, our method is significantly more efficient compared with other methods. Firstly, our approach swiftly learns to edit new attributes with a small training time (4 to 8 minutes), a notable contrast to PREIM3D’s requirement of several hours of training alongside the necessity of a pre-trained attribute-specific classifier. Secondly, despite employing a language encoder, our method’s inference time is comparable to that of PREIM3D. This efficiency stems from the fixed nature of the prompt post-training, allowing us to compute prompt features once and reuse them efficiently. Finally, our method introduces a sharing mechanism for style mappers across multiple attributes, delivering benefits in both storage and computation.
@@ 4 Experiment
Implementation Details. For the bulk of our experiments, we used a pre-trained 3D generative model GMPI [51] as our base 3D GAN and kept it frozen except for the alpha maps branch. To learn the attribute-specific prompt, we use a pre-trained CLIP [33] model. CLIP’s text encoder is utilized for both training and inference, and its image encoder is used during training only. To train our LAE module, we set $\lambda_{dclip}=1.0$, $\lambda_{sc}=0.8$, $\lambda_{id}=0.8$, $\lambda_{idvc}=0.5$, $\lambda_{latent}=0.5$ and $\lambda_{\alpha}=0.5$. We optimize using Adam [23] with a learning rate of $0.001$ and parameters $\beta_{1}=0.9$ and $\beta_{2}=0.95$. Similar to GMPI [51], we use $32$ planes for training and $96$ planes for inference and set near and far depth for MPI as $0.95/1.12$, use the depth normalization as [51]. To further evaluate, we integrate our method to other 3D generative model like EG3D [7], StyleNeRF [14] and CIP3D [52]. Our method is extremely efficient in terms of both computation and space requirements, as it only requires style tokens and style mappers to be trained and stored.
@@ 4.1 Quantitative Results
As GMPI [51] and EG3D [7] lacks attribute editing capabilities, we present in Table 3 a single averaged value across 1000 images. In contrast, our method calculates this value individually across three different attributes (smile, makeup, and age) using 1000 images for each attribute, reporting the average value. In a comparison of the depth and pose accuracy between GMPI [51], EG3D [7] and our method. As depicted in the table, our method faithfully preserves depth and pose accuracy across attributes. For instance, concerning smile attribute editing, our method exhibits a depth accuracy of 0.51, marginally lower than GMPI’s 0.49 (lower is better) while using GMPI backbone, when integrating the LAE Module with the EG3D we can observe a similar pattern. Similar observations are evident for makeup and age attributes. These findings highlight our method’s capability to maintain 3D geometry while facilitating attribute editing.
@@ 4.2 Qualitative Results
Firstly, to demonstrate the efficacy of our method, we showcase qualitative results and compare them with the GMPI and the state-of-the-art PREIM3D [25] in Figure 3. Following the approach outlined in [25], we uniformly sampled 9 images with yaw angles ranging from -30° to 30° and pitch angles between -20° and 20°. To facilitate comparative analysis, we utilize pre-trained weights provided by GMPI [51] and PREIM3D [25] and replicate their results.
@@
To showcase the efficacy of our method in preserving both identity and camera poses, we conduct a comparison with GMPI in Figure 4. This comparison spans two attributes observed across four randomly selected camera angles. The figure demonstrates how our method handles the editing of intricate and challenging attributes, such as specific emojis or blue eye colors, without compromising camera angles or altering the subject’s identity. Further, we use CelebA-HQ [19] to illustrate our method’s ability to edit real images. We first invert images to obtain their latent code using E4E [43]. Subsequently, these images are edited using our method, and the results of this experiment are shown in Figure 4(b). Our method maintains the identity and preserves the camera pose while successfully altering attributes. It’s essential to note that the preservation quality of our method is inherently constrained by GMPI.
@@ 4.3.2 Robustness and biasness
To delve deeper into understanding the influence of attributes on each other, such as how changing "blue eyes" may affect "skin tone" as shown in row 2 of Figure. 4(a), we explored the bias present in the CLIP model. Our method finds semantic direction in the latent space of 3D-aware GANs via. text, encoded by CLIP. Attribute editing accuracy and control rely on CLIP’s encoding, potentially inheriting biases, noted in previous works[3]. To understand the biases of the clip, we performed a simple experiment by encoding ten variations of prompts for white, brown, and black faces and blue eyes with CLIP and calculating the distance of the faces with blue eyes. We found that the encoding of the white face is closest to the blue eyes: white $0.035$, brown $0.051$, and black $0.065$.
4.4 Ablation Study
In Figure 7(a), highlights the impact of introducing proposed losses and the LAE Module into the 3D generation method for the attribute "curly hair". Integrating the LAE Module with $L_{LAE}$ loss brings about changes in the desired attribute, emphasizing the significance of using $L_{LAE}$ loss during the training of the LAE Module. However, despite altering the desired attributes, the LAE module falls short of preserving the identity and poses of the edited images. This issue is addressed by introducing $L_{id}+L_{idvc}$, resulting in images with desired attributes while maintaining a similar identity to the original image, though poses are still not adequately preserved. Our final approach, employing $L_{total}$, effectively preserves poses, identity, and generates 3D-aware edited images
@@ 4.4 Ablation Study
4.5 Conclusion
In this paper, we present an efficient pipeline for editing 3D-aware and view-consistent facial image attributes specified through textual prompts. Our method comprises a text-driven Latent Attribute Editor (LAE) alongside a 3D GAN. The LAE integrates learned style tokens and a style mapper, leveraging a pre-trained CLIP model to find semantic editing directions within 3D GAN latent space. The text-driven approach and utilization of style tokens render our method exceptionally efficient in handling novel editing directions. Additionally, we employ several 3D-aware identities and pose preservation losses to ensure view consistency in the generated images. Our method’s effectiveness is validated through a comprehensive array of qualitative and quantitative experiments.